A person pulls on a spring, stretching it 20 cm, which requires a maximum force of


800 N. Find the potential energy stored in the spring when it is stretched 20 cm and


compressed 5 cm.

1
Expert's answer
2022-02-09T06:26:19-0500

Given:

"x_1=\rm 20\: cm=0.2\: m"

"x_2=\rm 5\: cm=0.05\: m"

"F=\rm 800\: N"


The sping constant

"k=\frac{F}{x_1}=\frac{800\:\rm N}{0.2\:\rm m}=4000\:\rm N/m"

The potential energy stored in the spring:

(a)

"E_p=\frac{kx_1^2}{2}=\rm \frac{4000*0.2^2}{2}=80\: J"

(b)

"E_p=\frac{kx_2^2}{2}=\rm \frac{4000*0.05^2}{2}=5\: J"